The weekly sales of a store follow the normal distribution with a mean of 3000 euros and standard deviation 350 euros. What is the probability that the weekly sales are higher than 2800 euros?
Answer #1

X N (mean = , variance = o)

So,

Z-A-MN(0,1)

Thus,

(2) = P(Z < 2) = cdf of Z   ( we have to use standard normal table or software to find this value, we will use R Studio here, code: pnorm(z)

X= Weekly sales

X ~ N(mean = 3000, variance = 350)

P(x > 2800) = 1 - P(X < 2800) = 1 - P(4350 px - 3000 2800 - 3000 < 200,350 * = 1 – P(Z <- =1 – 11 = 1-0.2838546 = 0.7161454

R Studio code and output:

> pnorm(-4/7)
[1] 0.2838546